I think I know how to do these problems just need someone to double check my answers with

1 For the two vectors in the x-y plane, a) Calculate the sum of the vectors (R) by first calculating Rx and Ry (the scalar sums of the x and b) c) y components, respectively) and then writing the vector result R in ijk summation format. Draw the graphical representation of the summation of the two vectors, this should include x-y reference axes, the two original vectors, and the resultant R. Calculate the magnitude of the resultant, Calculate φ , the angle the resultant makes with the x-axis. d) 40° 352.5
Answer #1

There are two vectors in the figure. Let's say the first vector as overrightarrow{A} (having magnitude 3) and the other vector as overrightarrow{B} (having magnitude 2.5).

40° 35へ2.5

From the figure, we can write

overrightarrow{A} = 3 , cos , 40^{circ} , hat{i} + 3 , sin , 40^{circ} , hat{j}

or,   = 2.3i + 1.9

And, B = 2. 5 cos 350 (-J) +2.5 sin 35%

or,   B = 2.5sin 35%-2.5 cos 35°j

or, B = 1.43 i-204j

(a) The sum of the vectors :

  overrightarrow{R} = overrightarrow{A}+overrightarrow{B}

  ( 2.3 i + 1.95) + ( 1.43 i-204j

  = left (2.3 + 1.43 ight ) , hat{i}+ left ( 1.9 - 2.04 ight ) , hat{j}

or,   R-3.73-0.61

[ Here, Rx = 3.73 and Ry = - 0.61 ]

(b)  The graphical representation can be done by using shifting of vectors.

  R=A+B 2.5 3 40°

This is known as triangular rule of vector addition . You can also draw it using trapezoidal rule.

(c) Magnitude of vector R Can be calculated as

  R = left | overrightarrow{R} ight | = sqrt{R_{x}^{2} + R_{y}^{2}}

or,  R = sqrt{(3.73)^{2} + (-0.61)^{2}}

or,  R = sqrt{13.91 + 0.37}

or,  R = sqrt{14.285} = 3.8

(d) The angle ( heta) (or phi _{x} ) made by resultant vector with x - axis can be calculated as

  heta = tan ^{-1} left ( rac{R_{y}}{R_{x}} ight )

or, θ = tan-1 (-0.61 3.73

or,  heta = tan ^{-1} left ( -0.164 ight ) = -9.28 ^{circ}

Hence, the resultant vector makes -9.280 with positive x - axis or 350.70 (=360 - 9.28) with x - axis in anticlockwise direction.

9.2

For any doubt please comment and please give an up vote. Thank you.